Calculate the following information for Rudolphs Snowmobile Rentals Ine Total fixed costs Unit sales price Unit variable cost $7,800 $200 per hour S 50 per hour a) Compute the contribution margin per rental hour. Sales Vaviabit cost b) Compute the contribution margin ratio. c) Compute the break-even point in hours. d) Compute the break-even point in sales dollars. e) Calculate the hours needed to earn S29,700 in operating income.
ie sales price to $180 and has no changes itn vaniable or fixod conts, calculate the hoars necdod to becak even g) If Rudlphs decreascs the sakes price to $180 and has ao changes in variable or fixed costs, cakculate the hours needed to cars $29,700 in eperating income hi Supposc Rudolphs decides so pay their salcspeople os coemmission. With sales price at $200 per unit, variahlc costs per unit will icrcase by $20, while fixed costs will decrease by $2.600. What is the new break-even point in hoars 1) Suppose Kadolph s decides to pay their salespeople on commissioe. With sales price at S150 per unit, vanable costs per unit will increase by 515, whtle fixed of sental hours beoded to carm $29,700 in operating income? costs will decrease by $2,760. What is the number 5) Using the original data, peepare a ion margin format income statement for sales of 400 rental hoars
Answer #1

Total fixed cost = $7,800

Selling price = $200 per hour

Variable cost = $50 per hour

(a)

Contribution margin = Selling price - Variable cost

= 200 - 50

= $150 per hour

(b)

Contribution margin ratio = Contribution margin/Selling price

= 150/200

= 75%

(c)

Break even point = Fixed cost/Contribution margin

= 7,800/150

= 52 hours

(d)

Break even point ($) = Fixed cost/Contribution margin ratio

= 7,800/75%

= $10,400

(e)

Units to be sold to get a target profit = (Fixed cost + Target profit)/Contribution margin

Hence, hours needed to get an operating income of $29,700 = (7,800 + 29,700)/150

= 250 hours

(f)

Contribution margin = Selling price - Variable cost

= 180 - 50

= $130 per hour

Break even point = Fixed cost/Contribution margin

= 7,800/130

= 60 hours

(g)

Units to be sold to get a target profit = (Fixed cost + Target profit)/Contribution margin

Hence, hours needed to get an operating income of $29,700 = (7,800 + 29,700)/130

= 288.46 hours

Note: As per HOMEWORKLIB POLICY, I was required to answer first 4 parts but I have answered 7 parts. You please post your other queries separately. Thanks.

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.